A human trait that is caused by an autosomal dominant mutant allele has a penetrance of 50%. A man who is heterozygous for th6
Answer #1

50% percent penetrance means that half of the populations with the heterozygous condition will get the disease even if the mutation is dominant.

Let's make a punnet square with a father of genotype Aa and Mother, aa.

a a
A Aa Aa
a aa aa

From the above cross

The probability of showing normal phenotype = 3/4 (2 from aa and 1/2 from Aa as 50% penetrance is there)

The probability of having two children with normal phenotype = 3/4*3/4= 9/16